OpenStudy (anonymous):

how do I find the rate r=xp out of my outcome of p=-x+62 when I used the points 42,20 do I just multiply them together? to find r

OpenStudy (anonymous):

right......if u want to find xp=r using the point(42,20), then just multiply them

OpenStudy (amistre64):

if that is how r is defined; then yes; r = x * p

OpenStudy (anonymous):

I just dont think it is right here is the question A company’s revenue is the amount of money that comes in from sales, before business costs are subtracted. For a single product, you can find the revenue by multiplying the quantity of the product sold, x, by the demand equation, p. b. Substitute the result you found from part a. into the equation R = xp to find the revenue equation. Provide your answer in simplified form.

OpenStudy (anonymous):

what is part a ?

OpenStudy (anonymous):

the answer was p=-x+62

OpenStudy (anonymous):

this is the equation of line

OpenStudy (anonymous):

this is result of a 2 points in the y=mx+b slope

OpenStudy (anonymous):

sin ce the result was for p would I subsitute that in the r= xp ? making it r= x(-x+62)

OpenStudy (anonymous):

r= x^2-62x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so this is the rate?

OpenStudy (anonymous):

depends on x

OpenStudy (anonymous):

ok so 42 and 20

OpenStudy (anonymous):

42 is x

OpenStudy (anonymous):

but it is the point on the line

OpenStudy (anonymous):

the question just asks to use the results from part a so I figured?

OpenStudy (anonymous):

in part a u found p=-x+62, right? in part b u r asked to use part a to find r..right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and r=xp given.right?

OpenStudy (anonymous):

so r=x^2-64x is the answer using part a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or m not understanding what u want to solve :)

OpenStudy (anonymous):

so the equation is just the answer lol I was expecting to have to solve for the equation lol

OpenStudy (anonymous):

look it depends on the sale, x

OpenStudy (anonymous):

it totally says to find the revenue equation too sorry for the confustion

OpenStudy (anonymous):

confusion

OpenStudy (anonymous):

oh :)

OpenStudy (amistre64):

revenue (money coming in) = price * number sold

OpenStudy (anonymous):

thanks GURU G :P

OpenStudy (amistre64):

since price = (-x+62); you would multiply that by x to get the revenue

OpenStudy (amistre64):

:) r = -x^2 + 62x would be the revenue equation

OpenStudy (anonymous):

correct :)

OpenStudy (amistre64):

since r i snow in the fashion of an inverted parabola; you can determine the next question, which I assume is when do we have maximum profits :)